# The numeraire e-variable

### 29 Feb 2024 v1

@@ 6 Beyond the logarithm: reverse Rényi projection
for $\gamma>1$. In this section we outline the resulting theory using a well-known method from mathematical finance pioneered by Karatzas et al. (1991); Kramkov and Schachermayer (1999). However, in contrast to that literature we work as in Sections 2 and 3 with completely arbitrary ${\mathcal{P}}$ and ${\mathbb{Q}}$.
@@
$\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathbb{Q}}[U(X)].$
We denote the maximizer by $X^{*}_{\gamma}$. Observe that
${\mathbb{E}}_{\mathbb{Q}}[U^{\prime}(X^{*}_{\gamma})X^{*}_{\gamma}]={\mathbb{E}}_{\mathbb{Q}}\left[(X^{*}_{\gamma})^{1-\gamma}\right]=(1-\gamma){\mathbb{E}}_{\mathbb{Q}}[U(X^{*}_{\gamma})]\leq(1-\gamma)U(1)=1,$
where the inequality uses that the constant e-variable $X=1$ is suboptimal and that $\gamma>1$. Thus Lemma 2.8 further implies the first-order condition
${\mathbb{E}}_{\mathbb{Q}}\left[(X^{*}_{\gamma})^{-\gamma}X\right]\leq{\mathbb{E}}_{\mathbb{Q}}\left[(X^{*}_{\gamma})^{1-\gamma}\right],\quad X\in{\mathcal{E}},$ (6.1)
where $(X^{*}_{\gamma})^{-\gamma}X$ is to be understood as zero on $\{X^{*}_{\gamma}=\infty\}$. This will allow us to identify an analog of the RIPr which we denote by ${\mathbb{P}}^{*}_{\gamma}$. Let us first address the fully degenerate case where $X^{*}_{\gamma}=\infty$, ${\mathbb{Q}}$-almost surely, meaning that ${\mathbb{Q}}$ is concentrated on a ${\mathcal{P}}$-negligible set. In this case we simply set ${\mathbb{P}}^{*}_{\gamma}=0$. In the more interesting case where $X^{*}_{\gamma}$ is finite with positive ${\mathbb{Q}}$-probability, we let ${\mathbb{P}}^{*}_{\gamma}$ be the measure whose density with respect to ${\mathbb{Q}}$ is
$\frac{d{\mathbb{P}}^{*}_{\gamma}}{d{\mathbb{Q}}}=\frac{(X^{*}_{\gamma})^{-\gamma}}{{\mathbb{E}}_{\mathbb{Q}}\left[(X^{*}_{\gamma})^{1-\gamma}\right]}.$ (6.2)
Thanks to (6.1), ${\mathbb{P}}^{*}_{\gamma}$ belongs to ${\mathcal{P}}^{\circ\circ}$. In analogy with (3.5) we expect ${\mathbb{P}}^{*}_{\gamma}$ to minimize a suitable dual objective function. To discover its form, we look for a weak duality inequality and introduce the Legendre transform $V(y)=\sup_{x>0}\{U(x)-xy\}$ of $U$. Evaluating the supremum shows that
@@
Now, for any e-variable $X$, any ${\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}$, and any $y>0$, we have
${\mathbb{E}}_{\mathbb{Q}}[U(X)]$ $\leq{\mathbb{E}}_{\mathbb{Q}}[U(X)]+y(1-{\mathbb{E}}_{\mathbb{P}}[X])$
$\leq{\mathbb{E}}_{\mathbb{Q}}\left[U(X)-y\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}X\right]+y$
$\leq{\mathbb{E}}_{\mathbb{Q}}\left[V\left(y\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right)\right]+y$
$=-\frac{y^{1-1/\gamma}}{1-1/\gamma}{\mathbb{E}}_{\mathbb{Q}}\left[\left(\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right)^{1-1/\gamma}\right]+y.$
To obtain a tight bound we minimize the right-hand side over $y$. A computation shows that the minimizer is $y={\mathbb{E}}_{\mathbb{Q}}[(d{\mathbb{P}}^{a}/d{\mathbb{Q}})^{1-1/\gamma}]^{\gamma}$, which is finite, possibly zero, thanks to Jensen’s inequality. With this choice we get, for any e-variable $X$ and any ${\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}$, the weak duality inequality
${\mathbb{E}}_{\mathbb{Q}}[U(X)]\leq\frac{1}{1-\gamma}{\mathbb{E}}_{\mathbb{Q}}\left[\left(\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right)^{1-1/\gamma}\right]^{\gamma}.$ (6.3)
Equality is achieved by taking $X=X^{*}_{\gamma}$ and ${\mathbb{P}}={\mathbb{P}}^{*}_{\gamma}$ so we conclude that ${\mathbb{P}}^{*}_{\gamma}$ solves the minimization problem
$\inf_{{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}}\frac{1}{1-\gamma}{\mathbb{E}}_{\mathbb{Q}}\left[\left(\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right)^{1-1/\gamma}\right]^{\gamma}$
with optimal value equal to ${\mathbb{E}}_{\mathbb{Q}}[U(X^{*}_{\gamma})]$. Observe that the quantity being minimized is an increasing function of the Rényi divergence of order $1/\gamma$ of ${\mathbb{Q}}$ from ${\mathbb{P}}$,
$D_{1/\gamma}({\mathbb{Q}}\mid{\mathbb{P}})=\frac{1}{1/\gamma-1}\log{\mathbb{E}}_{\mathbb{Q}}\left[\left(\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right)^{1-1/\gamma}\right].$
Thus ${\mathbb{P}}^{*}_{\gamma}$ also minimizes the Rényi divergence, and as such it is known as the reverse Rényi projection of order $1/\gamma$ of ${\mathbb{Q}}$ on ${\mathcal{P}}^{\circ\circ}$. In particular, we see that the reverse Rényi projection always exists if the order $1/\gamma$ is in $(0,1)$. We summarize the above observations, along with a few others, in the following theorem.
@@
$\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathbb{Q}}[U(X)]=\frac{1}{1-\gamma}\exp\left((\gamma^{-1}-1)\inf_{{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}}D_{1/\gamma}({\mathbb{Q}}\mid{\mathbb{P}})\right),$ (6.4)
and the supremum and infimum are attained by $X^{*}_{\gamma}\in{\mathcal{E}}$ and ${\mathbb{P}}^{*}_{\gamma}\in{\mathcal{P}}^{\circ\circ}$ related by (6.2). Furthermore, $X^{*}_{\gamma}$ is the ${\mathbb{Q}}$-almost surely unique maximizer, and ${\mathbb{P}}^{*}_{\gamma}$ is the unique minimizer among the elements of ${\mathcal{P}}^{\circ\circ}$ that are absolutely continuous with respect to ${\mathbb{Q}}$. Finally, $X^{*}_{\gamma}$ can be expressed in terms of ${\mathbb{P}}^{*}_{\gamma}$ by
$X^{*}_{\gamma}=\left(\frac{d{\mathbb{P}}^{*}_{\gamma}}{d{\mathbb{Q}}}\right)^{-1/\gamma}\biggm{/}{\mathbb{E}}_{\mathbb{Q}}\left[\left(\frac{d{\mathbb{P}}^{*}_{\gamma}}{d{\mathbb{Q}}}\right)^{1-1/\gamma}\right],$ (6.5)
which is understood as +$\infty$+ ∞ on the set where $d{\mathbb{P}}^{*}_{\gamma}/d{\mathbb{Q}}=0$.
@@
The equality in (6.4) and the form of the solutions were shown above. The ${\mathbb{Q}}$-almost sure uniqueness of $X^{*}_{\gamma}$ follows from strict concavity of the function $f(X)={\mathbb{E}}_{\mathbb{Q}}[U(X)]$ on ${\mathcal{E}}$. Since $\gamma>1$, the Rényi divergence is a decreasing transformation of the function $g({\mathbb{P}})={\mathbb{E}}_{\mathbb{Q}}[(d{\mathbb{P}}^{a}/d{\mathbb{Q}})^{1-1/\gamma}]$, which is thus maximized by ${\mathbb{P}}^{*}_{\gamma}$. The function $g({\mathbb{P}})$ is strictly concave on the set $({\mathcal{P}}^{\circ\circ})^{a}$ of elements of ${\mathcal{P}}^{\circ\circ}$ that are absolutely continuous with respect to ${\mathbb{Q}}$. Thus ${\mathbb{P}}^{*}_{\gamma}$ is the unique maximizer of $g({\mathbb{P}})$, and the unique minimizer of the Rényi divergence. Finally, (6.5) is simply obtained by inverting (6.2). ∎
@@ 7 Summary
We established that under no assumptions on the composite null ${\mathcal{P}}$ and point alternative ${\mathbb{Q}}$, there exists a unique special e-variable $X^{*}$, the numeraire, which satisfies ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]\leq 1$ for every other e-variable $X$, and is also log-optimal. Further, $X^{*}$ can be written as a likelihood ratio $d{\mathbb{Q}}/d{\mathbb{P}}^{*}$ for a unique (sub-)probability distribution ${\mathbb{P}}^{*}$, absolutely continuous with respect to ${\mathbb{Q}}$ and belonging to the bipolar ${\mathcal{P}}^{\circ\circ}$ of ${\mathcal{P}}$ (the effective null hypothesis). The numeraire is the only e-variable of this form. A strong duality theory establishes that ${\mathbb{E}}_{\mathbb{Q}}[\log X^{*}]$ equals the minimum relative entropy of ${\mathbb{Q}}$ to ${\mathcal{P}}^{\circ\circ}$, achieved by ${\mathbb{P}}^{*}$. This fully generalizes (to composite nulls) the fact that the likelihood ratio is the log-optimal e-variable for point nulls. We gave several sufficient conditions to identify and certify the numeraire, and showed that these were easy to verify in several nonparametric examples without a reference measure that were out of the reach of earlier methods. Finally, we also showed how to generalize our theory beyond log-optimality.
Despite our focus on point alternatives, there are two natural ways in which this theory has immediate implications for composite alternatives ${\mathcal{Q}}$. The literature deals with composite alternatives in two ways (Ramdas et al., 2023). The first, called the ‘method of mixtures’, chooses a distribution over ${\mathcal{Q}}$, and the second, called the ‘plug-in method’, chooses a particular representative element ${\mathbb{Q}}$ of ${\mathcal{Q}}$ (this element is usually chosen by sample-splitting). In either case, one can apply our theory with the single mixture or plug-in distribution serving as our point alternative. In particular, our results imply that the relatively general method of universal inference (Wasserman et al., 2020) is inadmissible for testing composite ${\mathcal{P}}$ in all nondegenerate cases. In fact, our method is ‘truly universal’ in the sense that it imposes even weaker assumptions than their work (which needs a reference measure) and our numeraire always dominates their e-variable which uses the maximum likelihood in place of the RIPr.
Finally, our work has implications for sequential testing using nonnegative supermartingales and ‘e-processes’ (sequential generalizations of e-variables), a subject that introduces additional objects such as filtrations and stopping times. This is a direction we are currently pursuing.

### 22 Mar 2024 v2

@@ 6 Beyond the logarithm: reverse Rényi projection
for $\gamma>1$. In this section we outline the resulting theory using a well-known method from mathematical finance pioneered by Karatzas et al. (1991) and Kramkov and Schachermayer (1999). However, in contrast to that literature we work as in Sections 2 and 3 with completely arbitrary ${\mathcal{P}}$ and ${\mathsf{Q}}$.
@@
$\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[U(X)].$
We denote the maximizer by $X^{*}_{\gamma}$ and call it the $U$-optimal e-variable. Observe that
${\mathbb{E}}_{\mathsf{Q}}[U^{\prime}(X^{*}_{\gamma})X^{*}_{\gamma}]={\mathbb{E}}_{\mathsf{Q}}\left[(X^{*}_{\gamma})^{1-\gamma}\right]=(1-\gamma){\mathbb{E}}_{\mathsf{Q}}[U(X^{*}_{\gamma})]\leq(1-\gamma)U(1)=1,$
where we use the convention $U^{\prime}(\infty)\infty=0$ and where the inequality relies on the fact that the constant e-variable $X=1$ is suboptimal and that $\gamma>1$. Thus Lemma 2.8 further implies the first-order condition
${\mathbb{E}}_{\mathsf{Q}}\left[(X^{*}_{\gamma})^{-\gamma}X\right]\leq{\mathbb{E}}_{\mathsf{Q}}\left[(X^{*}_{\gamma})^{1-\gamma}\right],\quad X\in{\mathcal{E}},$ (6.1)
where $(X^{*}_{\gamma})^{-\gamma}X$ is understood as zero on $\{X^{*}_{\gamma}=\infty\}$. This will allow us to identify an analog of the RIPr which we denote by ${\mathsf{P}}^{*}_{\gamma}$. In the fully degenerate case where $X^{*}_{\gamma}=\infty$, ${\mathsf{Q}}$-almost surely, we simply set ${\mathsf{P}}^{*}_{\gamma}=0$. In the more interesting case where $X^{*}_{\gamma}$ is finite with positive ${\mathsf{Q}}$-probability, we let ${\mathsf{P}}^{*}_{\gamma}$ be the measure whose density with respect to ${\mathsf{Q}}$ is
$\frac{d{\mathsf{P}}^{*}_{\gamma}}{d{\mathsf{Q}}}=\frac{(X^{*}_{\gamma})^{-\gamma}}{{\mathbb{E}}_{\mathsf{Q}}\left[(X^{*}_{\gamma})^{1-\gamma}\right]}.$ (6.2)
Thanks to (6.1), ${\mathsf{P}}^{*}_{\gamma}$ belongs to ${\mathcal{P}}^{\circ\circ}$. In analogy with (3.3) we expect ${\mathsf{P}}^{*}_{\gamma}$ to minimize a suitable dual objective function. To discover its form, we look for a weak duality inequality and introduce the Legendre transform $V(y)=\sup_{x>0}\{U(x)-xy\}$. Evaluating the supremum shows that
@@
Now, for any e-variable $X$, any ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$, and any $y\geq 0$, we have
${\mathbb{E}}_{\mathsf{Q}}[U(X)]$ $\leq{\mathbb{E}}_{\mathsf{Q}}[U(X)]+y(1-{\mathbb{E}}_{\mathsf{P}}[X])$
$\leq{\mathbb{E}}_{\mathsf{Q}}\left[U(X)-y\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}X\right]+y$
$\leq{\mathbb{E}}_{\mathsf{Q}}\left[V\left(y\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right)\right]+y$
$=-\frac{y^{1-1/\gamma}}{1-1/\gamma}{\mathbb{E}}_{\mathsf{Q}}\left[\left(\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right)^{1-1/\gamma}\right]+y.$
To obtain a tight bound we minimize the right-hand side over $y$. A computation shows that the minimizer is $y={\mathbb{E}}_{\mathsf{Q}}[(d{\mathsf{P}}^{a}/d{\mathsf{Q}})^{1-1/\gamma}]^{\gamma}$, which is finite, possibly zero, thanks to Jensen’s inequality. With this choice we get, for any e-variable $X$ and any ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$, the weak duality inequality
${\mathbb{E}}_{\mathsf{Q}}[U(X)]\leq\frac{1}{1-\gamma}{\mathbb{E}}_{\mathsf{Q}}\left[\left(\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right)^{1-1/\gamma}\right]^{\gamma}.$
Equality is achieved by taking $X=X^{*}_{\gamma}$ and ${\mathsf{P}}={\mathsf{P}}^{*}_{\gamma}$, so we conclude that ${\mathsf{P}}^{*}_{\gamma}$ solves the minimization problem
$\inf_{{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}}\frac{1}{1-\gamma}{\mathbb{E}}_{\mathsf{Q}}\left[\left(\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right)^{1-1/\gamma}\right]^{\gamma}$
with optimal value equal to ${\mathbb{E}}_{\mathsf{Q}}[U(X^{*}_{\gamma})]$. Observe that the quantity being minimized is an increasing function of the Rényi divergence of order $1/\gamma$ of ${\mathsf{Q}}$ from ${\mathsf{P}}$,
$D_{1/\gamma}({\mathsf{Q}}\mid{\mathsf{P}})=\frac{1}{1/\gamma-1}\log{\mathbb{E}}_{\mathsf{Q}}\left[\left(\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right)^{1-1/\gamma}\right].$
Thus ${\mathsf{P}}^{*}_{\gamma}$ also minimizes the Rényi divergence, and as such it is known as the reverse Rényi projection of order $1/\gamma$ of ${\mathsf{Q}}$ on ${\mathcal{P}}^{\circ\circ}$. In particular, we see that the reverse Rényi projection always exists if the order $1/\gamma$ is in $(0,1)$. We summarize the above observations, along with a few others, in the following theorem.
@@
$\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[U(X)]=\frac{1}{1-\gamma}\exp\left((\gamma^{-1}-1)\inf_{{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}}D_{1/\gamma}({\mathsf{Q}}\mid{\mathsf{P}})\right)^{\gamma},$ (6.3)
and the supremum and infimum are attained by $X^{*}_{\gamma}\in{\mathcal{E}}$ and ${\mathsf{P}}^{*}_{\gamma}\in{\mathcal{P}}^{\circ\circ}$ related by (6.2). Furthermore, $X^{*}_{\gamma}$ is the ${\mathsf{Q}}$-almost surely unique maximizer, and ${\mathsf{P}}^{*}_{\gamma}$ is the unique minimizer among the elements of ${\mathcal{P}}^{\circ\circ}$ that are absolutely continuous with respect to ${\mathsf{Q}}$. Finally, $X^{*}_{\gamma}$ can be expressed in terms of ${\mathsf{P}}^{*}_{\gamma}$ by
$X^{*}_{\gamma}=\left(\frac{d{\mathsf{P}}^{*}_{\gamma}}{d{\mathsf{Q}}}\right)^{-1/\gamma}\biggm{/}{\mathbb{E}}_{\mathsf{Q}}\left[\left(\frac{d{\mathsf{P}}^{*}_{\gamma}}{d{\mathsf{Q}}}\right)^{1-1/\gamma}\right],$ (6.4)
which is understood as +$\infty$+ ∞ on the set where $d{\mathsf{P}}^{*}_{\gamma}/d{\mathsf{Q}}=0$.
@@
The equality in (6.3) and the form of the solutions were shown above. The ${\mathsf{Q}}$-almost sure uniqueness of $X^{*}_{\gamma}$ follows from strict concavity of the function $f(X)={\mathbb{E}}_{\mathsf{Q}}[U(X)]$ on ${\mathcal{E}}$. Since $\gamma>1$, the Rényi divergence is a decreasing transformation of the function $g({\mathsf{P}})={\mathbb{E}}_{\mathsf{Q}}[(d{\mathsf{P}}^{a}/d{\mathsf{Q}})^{1-1/\gamma}]$, which is thus maximized by ${\mathsf{P}}^{*}_{\gamma}$. The function $g({\mathsf{P}})$ is strictly concave on the set of elements of ${\mathcal{P}}^{\circ\circ}$ that are absolutely continuous with respect to ${\mathsf{Q}}$. Thus ${\mathsf{P}}^{*}_{\gamma}$ is the unique maximizer of $g({\mathsf{P}})$, and the unique minimizer of the Rényi divergence. Finally, (6.4) is obtained by noting that
${\mathbb{E}}_{\mathsf{Q}}\left[(X^{*}_{\gamma})^{1-\gamma}\right]={\mathbb{E}}_{\mathsf{Q}}\left[\left(\frac{d{\mathsf{P}}^{*}_{\gamma}}{d{\mathsf{Q}}}\right)^{1-1/\gamma}\right]^{\gamma}$
and inverting (6.2). ∎
Remark 6.2.
Lardy et al. (2023, Section V) mention the e-variable $(d{\mathsf{P}}^{*}_{\gamma}/d{\mathsf{Q}})^{-1/\gamma}$ constructed using the Rényi projection. This e-variable is strictly smaller than $X^{*}_{\gamma}$, because the denominator in (6.4) is strictly smaller than one by Jensen’s inequality.
7 Composite alternatives and universal inference
We now briefly discuss implications of our work to composite alternatives. We purposefully keep the discussion informal.
Method of mixtures and plug-in method.
Despite our focus on point alternatives, there are two natural ways in which this theory has immediate implications for composite alternatives ${\mathcal{Q}}$. When testing ${\mathcal{P}}$ against ${\mathcal{Q}}$ for i.i.d. data $(Z_{t})_{t\in{\mathbb{N}}}$, one sensible goal is to derive e-variables $X_{t}$ depending on the sample $Z_{1},\ldots,Z_{t}$ that are ‘asymptotically optimal’ in the sense that for all ${\mathsf{Q}}\in{\mathcal{Q}}$, we would have ${\mathbb{E}}_{{\mathsf{Q}}^{\otimes t}}[t^{-1}\log X_{t}]\to H({\mathsf{Q}}\mid{\mathsf{P}}^{*}({\mathsf{Q}}))$ as the sample size $t$ increases to infinity, where we use ${\mathsf{P}}^{*}({\mathsf{Q}})$ to denote the RIPr of ${\mathsf{Q}}$ onto ${\mathcal{P}}^{\circ\circ}$. Because $H({\mathsf{Q}}\mid{\mathsf{P}}^{*}({\mathsf{Q}}))$ is the highest achievable growth rate if the data is generated according to ${\mathsf{Q}}$, the above says that this optimal growth rate is achieved asymptotically no matter which ${\mathsf{Q}}\in{\mathcal{Q}}$ is the true data generating distribution. This is clearly a desirable feature whenever the data comes from an unknown element of ${\mathcal{Q}}$.
The literature deals with composite alternatives in two ways (Ramdas et al., 2023). The first is called the method of mixtures: using a distribution $\pi$ over ${\mathcal{Q}}$ and assuming for simplicity that the elements of ${\mathcal{Q}}$ are mutually equivalent, one takes the mixture likelihood ratio $X_{t}=\int_{\mathcal{Q}}\prod_{i\leq t}\frac{d{\mathsf{Q}}}{d{\mathsf{P}}^{*}({\mathsf{Q}})}(Z_{i})\pi(d{\mathsf{Q}})$. The second method is called the plug-in method: using part of the sample $Z_{1},\ldots,Z_{t}$ one chooses a particular representative element $\hat{\mathsf{Q}}_{t}$ of ${\mathcal{Q}}$ and uses this to obtain the corresponding numeraire as a function of the remaining part of the sample. This method is an instance of sample-splitting, used for instance in universal inference (Wasserman et al., 2020). In either case we get valid e-variables, and under relatively weak conditions on $\pi$ or $\hat{\mathsf{Q}}_{t}$ one can prove that the resulting e-variable is indeed asymptotically optimal. Establishing tight necessary and sufficient conditions for such asymptotic optimality is left to future work.
Comparison to universal inference.
We assume that there exists a common reference measure $\mu$, and follow the notational conventions introduced before Theorem 4.7 to identify distributions with their respective densities (written in lowercase). For a point alternative $q$ over the data $Z$, the method of universal inference (Wasserman et al., 2020) boils down to constructing the e-variable $X^{\text{UI}}=q(Z)/p_{\text{max}}(Z)$ where $p_{\text{max}}(Z)=\sup_{p\in{\mathcal{P}}}p(Z)$ is the maximum likelihood. (Strictly speaking, the supremum here should be understood as an essential supremum under the reference measure.) To check that this is indeed an e-variable, simply note that for any ${\mathsf{P}}\in{\mathcal{P}}$ with density $p$ we have
${\mathbb{E}}_{\mathsf{P}}\left[\frac{q(Z)}{p_{\text{max}}(Z)}\right]\leq{\mathbb{E}}_{\mathsf{P}}\left[\frac{q(Z)}{p(Z)}\right]=1.$
To compare the universal inference e-variable with the numeraire $X^{*}=q(Z)/p^{*}(Z)$, we first claim that the RIPr satisfies $p^{*}(Z)\leq p_{\text{max}}(Z)$ up to $\mu$-nullsets. This is obvious if the RIPr belongs to ${\mathcal{P}}$, but in general it only belongs to ${\mathcal{P}}^{\circ\circ}$. In this case the claim follows by applying the following lemma with $f=p_{\text{max}}$.
Lemma 7.1.
Let $f$ be a random variable such that $f\geq p$, $\mu$-almost surely, for all $p\in{\mathcal{P}}$. Then $f\geq p$, $\mu$-almost surely, for all $p\in{\mathcal{P}}^{\circ\circ}$.
Proof.
It is clear that $f\geq p$, $\mu$-almost surely, for all sub-probability densities $p$ dominated by some element of the convex hull of ${\mathcal{P}}$, and then also for all $p$ in the $\mu$-probability closure of this set. This was called ${\mathcal{P}}^{\prime\prime}$ in the proof of Theorem 4.7, where it was shown that ${\mathcal{P}}^{\prime\prime}$ is closed, convex, and solid. By Lemma 4.5, ${\mathcal{P}}^{\prime\prime}$ contains ${\mathcal{P}}^{\circ\circ}$, completing the proof. ∎
We thus see that $p^{*}(Z)\leq p_{\text{max}}(Z)$, and hence $X^{*}\geq X^{\text{UI}}$, up to $\mu$-nullsets. In nondegenerate situations involving a composite null hypothesis, the inequality will be strict with positive $\mu$-probability. Thus, in such cases, the relatively general method of universal inference is in fact inadmissible.
We end by noting that our numeraire is ‘truly universal’ in the sense that it imposes even weaker assumptions than the universal inference work, which needs a reference measure to define likelihoods.
@@ 8 Summary
We established that under no assumptions on the composite null ${\mathcal{P}}$ and point alternative ${\mathsf{Q}}$, there exists a unique special e-variable $X^{*}$, the numeraire, which satisfies ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for every other e-variable $X$, and is also log-optimal. Further, $X^{*}$ can be written as a likelihood ratio $d{\mathsf{Q}}/d{\mathsf{P}}^{*}$ for a unique (sub-)probability distribution ${\mathsf{P}}^{*}$, absolutely continuous with respect to ${\mathsf{Q}}$ and belonging to the bipolar ${\mathcal{P}}^{\circ\circ}$ of ${\mathcal{P}}$ (the effective null hypothesis). The numeraire is the only e-variable of this form. A strong duality theory establishes that ${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]$ equals the minimum relative entropy of ${\mathsf{Q}}$ to ${\mathcal{P}}^{\circ\circ}$, achieved by ${\mathsf{P}}^{*}$. This fully generalizes (to composite nulls) the fact that the likelihood ratio is the log-optimal e-variable for point nulls. We gave several sufficient conditions to identify and certify the numeraire, and showed that these were easy to verify in several nonparametric examples without a reference measure that were out of the reach of earlier methods. We also showed how to generalize our theory beyond log-optimality. Our work has implications for composite alternatives and for sequential testing using nonnegative supermartingales and ‘e-processes’ (sequential generalizations of e-variables). The latter topic introduces additional subtleties through dealing with filtrations and stopping times, a direction we are currently pursuing.